# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v4

@@ 1 Introduction
Researchers aim to create universal robots that can seamlessly integrate into human life to boost productivity, much like those depicted in the movie ”I, Robot”. However, a key hurdle in achieving this lies in the robots’ ability to manipulate objects in their unstructured environments according to different tasks. There is abundant literature available for improving the general manipulation capability of robots, which can be roughly categorized into model-based and learning-based approaches (Zarrin et al. (2023)). The real world is too diverse for universal robots and they must adapt to unstructured environments and arbitrary objects to manipulate effectively. Therefore, learning-based methods are crucial for manipulation tasks (Kleeberger et al. (2020)).
@@
This paper is structured as follows: In Sec. 2, we present a comprehensive framework of robot learning for general manipulation, based on the developmental history of robot learning for manipulation and general manipulation definition. We elaborate on the impact of foundation models on each module in the framework in the following sections. Sec. 3 is Human/Agnet Interaction module, Sec. 4 is Pre- and Post-conditions Detection module, Sec. 5 is Hierarchy of Skills module, Sec. 6 is State Perception module, Sec. 7 is Policy module, Sec. 8 is Manipulation Data Generation module. In Sec. 9, we discuss several issues of particular concern to us. In Sec. 10, we summarize the contributions of this survey and identify the limitations of the current framework as well as the challenges in each module.
@@ 2 Framework of Robot Learning for General Manipulation
Figure 2: Framework of Robot Learning for General Manipulation. The Pre-conditions Detection module $P$ perceives the environment to identify objects and the affordances objects support. The Interaction module $I$ receives instruction from a human or other agent. It uses perception information from the Pre-conditions Detection module $P$ to check for ambiguities in the instruction. If there are any ambiguities, it generates a question to clarify the instruction by asking the human or other agent. The Hierarchy of Skills module $H$ generates subgoals by using precise instruction from the Interaction module $I$ and perception information from the Pre-conditions Detection module $P$. Each subgoal is then passed to the Skill Execution module. In the Skill Execution module, Policy module $\Pi$ generates Action $\alpha$ based on the State $S$. To obtain the next state after executing the current action, State $S$ can either perceive it from the environment or use the Transition module $T$. To train the Skill Execution module, including the State module $S$, the Policy module $\Pi$ and the Transition module $T$, the Manipulation Data Generation module is required. This module provides a task-level manipulation dataset. When issues arise during execution, corrective instruction is sent to the Policy module $\Pi$ for manual adjustment. Policy module $\Pi$ modifies the current action to corrective action and saves corrective demonstration to the dataset for self-improvement of Policy module $\Pi$. After skill execution, Post-conditions Detection module $P$ determines the success of execution. If successful, proceed to the next subgoal; if not, the failure reason is conveyed to Post-hoc Correction module for self-correction.
@@ 3 Human/Agent Interaction
When conveying task instruction to the robot, there may contain language ambiguity in the task goal, such as having both red and green cups in the scene, and the task instruction is ’grasp the cup.’ This ambiguity may confuse the robot regarding which color cup to grasp. To address this issue, the robot needs to inquire about and confirm the final task objective from the human or other agent, thus requiring enhancement of their capability in text generation and comprehension. When conveying corrective instruction to a robot, it needs to comprehend the meaning of the corrective instruction and translate corrective instruction into appropriate actions. If needed, the robot should be able to convey its current state to human or agent. For instance, if a robot is picking up a book from a shelf filled with books, lifting too quickly may cause other books to fall. Human or collaborating agent need to alert the robot that the current lifting action is dangerous and advise it to lift slowly. If necessary, the robot should also report its current execution state, such as its grasping speed, and inquire whether this speed is considered high. However, corrective instruction are diverse; thus, understanding them is essential.
@@
As for corrective instruction, LILAC (Cui et al. (2023)) utilizes GPT-3 to distinguish between task instruction and corrective instruction. It then employs Distil-RoBERTa to extract text features and input them into the network to modify the robot’s original trajectory. LATTE (Bucker et al. (2023)), on the other hand, employs BERT and CLIP to extract features from corrective instruction and observation images and input them into the network to modify the robot’s original trajectory.
@@ 4 Pre- and Post-conditions Detection
In pre- and post-conditions detection, it is necessary to identify the initial and termination conditions. In pre-conditions detection, recognize objects and observe the affordances of objects. In post-conditions detection, identify whether a task has been successfully executed and provide reasons for task failure after skill execution. Currently, there are few papers focusing on identifying termination conditions. Cui et al. (2022) utilizes CLIP to compare the target’s text or image with the termination environment to determine the success of task execution. Few articles are found in this study that address the output of task failure reasons after skill execution. RobotGPT (Jin et al. (2024)) analysis task failure utilizes the positions of manipulated objects after execution, but task failure should be determined during execution. Therefore, this section focuses on literature discussing foundation models in pre-conditions detection including object affordance and object recognition.
@@ 4.1 Object Affordance
The methods mentioned above utilize foundation models to assist other learning methods in improving affordance maps or grasp poses. There are also direct approaches using foundation models to generate affordance maps and grasp poses. PartSLIP (Liu et al. (2023d)) converts 3D point clouds into 2D rendering images and inputs multi-view 2D images and textual descriptions of object parts into GLIP for object parts detection, ultimately fusing 2D bounding boxes into 3D segmentation to generate affordance maps. LAN-grasp (Mirjalili et al. (2023)) inputs human instruction into LLM, utilizing its prior knowledge to output the shape of part to be grasped. These shapes, along with the object’s 2D image, are then inputted into VLM to detect the bounding box for the grasping part. Finally, the bounding box and the point cloud from object 3D reconstruction are inputted into the grasp planner to generate grasp poses.
@@ 4.2 Object Recognition
The method for open-set detection on 2D images can be extended to the research direction of open-set detection on 3D point clouds. PointCLIP (Zhang et al. (2022)) utilizes pre-trained CLIP to extract multi-view depth image features of point cloud, then compares the extracted features with textual features to identify the point cloud category.
@@
As shown in Fig. 4, LLM provides object part-level knowledge via text, aiding in affordance map or grasp pose generation. While foundation models can be used for affordance, their performance remains unstable. In object recognition, representation learning aligns multimodal features with text, improving model cognition, similar to human think with words. It also supports open-set perception tasks, like detection and segmentation.
@@ 5 Hierarchy of skills
In the last two decades, research on skill hierarchy has led to powerful domain-independent planners and various real-world applications (Silver et al. (2022)). Models for skill hierarchy can be trained using text or videos, similar to how humans learn assembly procedures from instructional manuals or tutorial videos. As for tutorial videos, VLaMP (Patel et al. (2023) and SeeDo (Wang et al. (2024a)) train models to understand human video operations and HourVideo (Chandrasegaran et al. (2024)) proposes a benchmark dataset specifically designed for hour-long video-language understanding. Currently, the text-based approach dominates the skill hierarchy domain. PDDL, a Lisp-like language (Silver et al. (2022)), is commonly used to address skill hierarchy issues. However, as LLMs excel in natural language tasks and PDDL is not a natural language, researchers are exploring how LLMs can be employed by robot for skill hierarchy tasks (Vemprala et al. (2023); Jansen (2020); Driess et al. (2023)). Additionally, various benchmarks such as PlanBench (Valmeekam et al. (2023)) and CALVIN (Mees et al. (2022)) can assess the planning and reasoning capability of LLMs.
LLMs possess a notable limitation: they lack practical experience, hindering their utility for decision-making within a specific context, so the output of LLMs often cannot be translated into executable actions for the robot. Huang et al. (2022) first use pre-trained causal LLM to break down high-level tasks into logical mid-level action plans. Then, a pre-trained masked LLM is employed to convert mid-level action plans into admissible actions. The pre-trained causal LLM and pre-trained masked LLM can be the same LLM model, which plays different roles depending on various prompts. However, prompts usually require the context of the robot’s capability, its current state, and the environment. LLMs are considered ”forgetful” and don’t treat information in the system prompt as absolute. Despite efforts to reinforce task constraints in the objective prompt and extract numerical task contexts from the system prompt, storing them in data structures, errors caused by LLM forgetfulness remain unresolved (Chen and Huang (2023)).
@@
Regarding user preferences, TidyBot (Wu et al. (2023b)) trains LLM by collecting users’ preference data, enabling the trained LLM to choose behaviors that better align with user preferences. As for safety regulations, Yang et al. (2023d) incorporate ISO 61508, a global standard for safely deploying robots in industrial factory settings, into the constraints of the action sequence generation. As for the skill library, BOSS (Zhang et al. (2023c)) suggests using LLMs’ rich knowledge to guide skills chaining in the skill library, aiming to create new skills through combinations. RoboGen (Wang et al. (2023e)) employs generative models to create new skill task scenarios, then utilizes either reinforcement learning or gradient optimization methods to automatically learn new skills based on the reward function generated by the LLM. As for action execution feedback, REACT (Yao et al. (2022b)), COWP (Ding et al. (2022)), LLM-Planner (Song et al. (2023)), CoPAL (Joublin et al. (2023)) provide feedback on robot action execution to LLMs. This allows LLMs to adjust action sequences based on execution status, creating a closed-loop process for generating action sequences.
@@
When utilizing LLMs for skill hierarchy, it provides generalization. Regardless of whether the prompting input to LLMs is in natural language or PDDL format (Silver et al. (2022)), the hierarchy of skills possessed by LLMs or VLMs still exhibits instability. Hence, researchers are exploring approaches that integrate LLMs with classical PDDL-based planning methods for the hierarchy of skills. LLM+P (Liu et al. (2023a)) utilize LLMs to translate natural language into PDDL and input into a classical planner for the hierarchy of skills. Xie et al. (2023b) indicate that LLMs exhibit greater efficacy in translation tasks as opposed to planning.
@@
Following Fig. 5, we focus on the application of video in hierarchy of skills. Video data, being internet-based and possessing large-scale attributes, enables robust generalization when learning high-level task operation descriptions mapped from videos. This approach not only aids in hierarchy of skills but also supports one-shot and few-shot learning from human video. By watching a person perform a task, the video is broken down into steps. Each step is then executed using the skill library. Utilizing LLMs for hierarchy of skills, we emphasize the GD (Huang et al. (2023d)) framework, which circumvents the high cost of LLM fine-tuning. LLMs exhibit powerful translation abilities, thus offering significant utility in assisting classical planner.
@@ 6 State
These consumer cameras typically use principles such as structured light, time of flight, and traditional photometric stereo for depth estimation (Zhou et al. (2024)) and 3D representation can be generated by registering them using camera poses (Huang et al. (2024b)). However, when surfaces are shiny, bright, transparent, textureless or distant from the camera, depth images produced by consumer cameras are often noisy and incomplete. Several studies have addressed this challenging problem by learning to restore depth images (Dai et al. (2022); Fang et al. (2022); Sajjan et al. (2020)). However, the correct depth information may already be lost in the original depth. ASGrasp (Shi et al. (2024)) demonstrates that 3D reconstruction using raw multi-view images from consumer cameras is better than restoring the original depth. Many studies use a single image for 3D reconstruction (Fu et al. (2021)). However, a single image loses a significant amount of information, resulting in lower accuracy. Despite this, the zero-shot capability of current single image 3D reconstruction has led to its widespread application in simulation scene generation (in Sec. 8.2).
The 3D representation for 3D reconstruction can be expressed as explicit and implicit expressions (Zhou et al. (2024)). Explicit expressions include point clouds (Shi et al. (2024)), voxels (Jiang et al. (2021)), and meshes (Wen et al. (2019)). The three representations can be converted into each other (Jiang et al. (2021)), but each has its own advantages. A point cloud is made up of discrete points in space, providing flexibility in processing. A voxel can store spatial information inside an object but comes with high space complexity. A mesh uses triangle meshes to represent complex shapes and details accurately, such as deformation (Wen et al. (2019)). It ensures the projection is always convex, making it easier to rasterize (Zhou et al. (2024)). Implicit expressions represent 3D geometry using a function, such as signed distance function (SDF) (Chabra et al. (2020)), occupancy field (Jiang et al. (2021)), and radiance field (Mildenhall et al. (2021)). They offer differentiability and efficient storage, making them a powerful tool. In contrast, explicit expressions tend to be more intuitive.
@@
The image features output by NeRF can be used to build a relevancy map. They can also be lifted into 3D space through multi-view pose, serving as 3D features for downstream tasks (Ze et al. (2023)). 3D-LLM (Hong et al. (2023b)) extracts 2D features from multi-view rendered images using the CLIP image encoder. These features are then fused into 3D features through Direct Reconstruct, gradSLAM (Jatavallabhula et al. (2023)), or Neural Field methods (Hong et al. (2023b)), endowing 3D features with semantic information.

### v5

@@ 1 Introduction
Researchers aim to create universal robots that can seamlessly integrate into human life to boost productivity, much like those depicted in the movie “I, Robot”. However, a key hurdle in achieving this lies in the robots’ ability to manipulate objects in their unstructured environments according to different tasks. There is abundant literature available for improving the general manipulation capability of robots, which can be roughly categorized into model-based and learning-based approaches (Zarrin et al. (2023)). The real world is too diverse for universal robots and they must adapt to unstructured environments and arbitrary objects to manipulate effectively. Therefore, learning-based methods are crucial for manipulation tasks (Kleeberger et al. (2020)).
@@
This paper is structured as follows: In Sec. 2, we present a comprehensive framework of robot learning for general manipulation, based on the developmental history of robot learning for manipulation and general manipulation definition. We elaborate on the impact of foundation models on each module in the framework in the following sections. Sec. 3 is Human/Agent Interaction module, Sec. 4 is Pre- and Post-conditions Detection module, Sec. 5 is Hierarchy of Skills module, Sec. 6 is State Perception module, Sec. 7 is Policy module, Sec. 8 is Manipulation Data Generation module. In Sec. 9, we discuss several issues of particular concern to us. In Sec. 10, we summarize the contributions of this survey and identify the limitations of the current framework as well as the challenges in each module.
@@ 2 Framework of Robot Learning for General Manipulation
Figure 2: Framework of Robot Learning for General Manipulation. The Pre-conditions Detection module $P$ perceives the environment to identify objects and the affordances objects support. The Interaction module II receives instruction from a human or other agent. It uses perception information from the Pre-conditions Detection module $P$ to check for ambiguities in the instruction. If there are any ambiguities, it generates a question to clarify the instruction by asking the human or other agent. The Hierarchy of Skills module $H$ generates subgoals by using precise instruction from the Interaction module $I$ and perception information from the Pre-conditions Detection module $P$. Each subgoal is then passed to the Skill Execution module. In the Skill Execution module, Policy module $\Pi$ generates Action $\alpha$ based on the State $S$. To obtain the next state after executing the current action, State $S$ can either perceive it from the environment or use the Transition module $T$. To train the Skill Execution module, including the State module $S$, the Policy module $\Pi$ and the Transition module $T$, the Manipulation Data Generation module is required. This module provides a task-level manipulation dataset. When issues arise during execution, corrective instruction is sent to the Policy module $\Pi$ for manual adjustment. Policy module $\Pi$ modifies the current action to corrective action and saves corrective demonstration to the dataset for self-improvement of Policy module $\Pi$. After skill execution, Post-conditions Detection module $P$ determines the success of execution. If successful, proceed to the next subgoal; if not, the failure reason is conveyed to Post-hoc Correction module for self-correction.
@@ 3 Human/Agent Interaction
When conveying task instruction to the robot, there may contain language ambiguity in the task goal, such as having both red and green cups in the scene, and the task instruction is ’grasp the cup.’ This ambiguity may confuse the robot regarding which color cup to grasp. To address this issue, the robot needs to inquire about and confirm the final task objective from the human or other agent, thus requiring enhancement of their capability in text generation and comprehension. When conveying corrective instruction to a robot, it needs to comprehend the meaning of the corrective instruction and translate corrective instruction into appropriate actions. For instance, if a robot is picking up a book from a shelf filled with books, lifting too quickly may cause other books to fall. Human or collaborating agent need to alert the robot that the current lifting action is dangerous and advise it to lift slowly. If necessary, the robot should also report its current execution state, such as its grasping speed, and inquire whether this speed is considered high. However, corrective instruction are diverse; thus, understanding them is essential.
@@
As for corrective instruction, LILAC (Cui et al. (2023)) utilizes GPT-3 to distinguish between task instruction and corrective instruction. It then employs Distil-RoBERTa to extract text features and input them into the network to modify the robot’s original trajectory. LATTE (Bucker et al. (2023)), on the other hand, employs BERT and CLIP to extract features from corrective instruction and observation images and input them into the network to modify the robot’s original trajectory. RT-H (Belkhale et al. (2024)) employs VLMs in a two-step operation, initially outputting abstract delta-pose representations like ”move left,” which are then converted into delta poses and human intervention can enable robots to adjust trajectories based on human language instruction.
@@ 4 Pre- and Post-conditions Detection
In pre- and post-conditions detection, it is necessary to identify the initial and termination conditions. In pre-conditions detection, recognize objects and observe the affordances of objects. In post-conditions detection, identify whether a task has been successfully executed and provide reasons for task failure after skill execution. Currently, there are few papers focusing on identifying termination conditions. Cui et al. (2022) utilizes CLIP to compare the target’s text or image with the termination environment to determine the success of task execution. Few articles are found in this study that address the output of task failure reasons after skill execution. RobotGPT (Jin et al. (2024)) analysis task failure utilizes the positions of manipulated objects after execution, but task failure should be determined during execution. AHA (Duan et al. (2024)) uses a large number of robotic failure trajectories to fine-tune the VLM. The fine-tuned VLM leverages keyframe trajectory images and task descriptions from the robot’s current task execution process to detect failures and provide detailed, adaptable failure explanations. Therefore, this section focuses on literature discussing foundation models in pre-conditions detection including object affordance and object recognition.
@@ 4.1 Object Affordance
The methods mentioned above utilize foundation models to assist other learning methods in improving affordance maps or grasp poses. There are also direct approaches using foundation models to generate affordance maps and grasp poses. PartSLIP (Liu et al. (2023d)) converts 3D point clouds into 2D rendering images and inputs multi-view 2D images and textual descriptions of object parts into GLIP for object parts detection, ultimately fusing 2D bounding boxes into 3D segmentation to generate affordance maps. PartSLIP requires manual definition text prompts and additional algorithms to convert 2D boxes back to 3D regions. UAD (Tang et al. (2025)) clusters object points into fine-grained semantic regions based on pixel-wise features extracted from multi-view rendered images using DINOv2. It then queries the VLM to generate a set of task instructions and associates these instructions with the most relevant clustered region to construct the affordance map. LAN-grasp (Mirjalili et al. (2023)) inputs human instruction into LLM, utilizing its prior knowledge to output the shape of part to be grasped. These shapes, along with the object’s 2D image, are then inputted into VLM to detect the bounding box for the grasping part. Finally, the bounding box and the point cloud from object 3D reconstruction are inputted into the grasp planner to generate grasp poses.
@@ 4.2 Object Recognition
The method for open-set detection on 2D images can be extended to the research direction of open-set detection on 3D point clouds. PointCLIP (Zhang et al. (2022b)) utilizes pre-trained CLIP to extract multi-view depth image features of point cloud, then compares the extracted features with textual features to identify the point cloud category.
@@
As shown in Fig. 4, LLM provides object part-level knowledge via text, aiding in affordance map or grasp pose generation. Reinforcement learning can make robotic systems perform better though interaction than supervised learning trained on datasets. Direct use of foundation models avoids training. However, stability remains a concern. In object recognition, representation learning aligns multimodal features with text, improving model cognition, similar to human think with words. It also supports open-set perception tasks, like detection and segmentation.
@@ 5 Hierarchy of skills
The skill hierarchy is closely related to the field of task and motion planning (TAMP). TAMP aims to address high-level instructions by organizing tasks in a sequence that ensures dynamic feasibility (Guo et al. (2023)). There are three main types of classical TAMP methods: constraint-based TAMP, sampling-based TAMP, and optimization-based TAMP (Zhao et al. (2024)). Constraint-based and sampling-based TAMP define the problem with goal conditions. Unlike optimization-based TAMP, these approaches often cannot assess or compare the quality of the generated plan or final state due to the lack of objective functions, such as when the goal is to pour as much water as possible into the cup (Zhao et al. (2024); Zhang et al. (2022a)). However, optimization-based TAMP is sensitive to the initial conditions and goal setup of the problem (Zhao et al. (2024)).
The scalability of classical TAMP methods is often constrained by the tree search problem size for complex tasks and the computational cost of evaluating heuristics and optimal trajectories (Zhao et al. (2024)). Integrating learning-based approaches into TAMP enables informed decision-making based on prior examples and experiences and improves flexibility and generalizability (Guo et al. (2023)). Models for skill hierarchy can be trained using text or videos, similar to how humans learn assembly procedures from instructional manuals or tutorial videos. As for tutorial videos, VLaMP (Patel et al. (2023) and SeeDo (Wang et al. (2024a)) use trained models to understand human video operations and HourVideo (Chandrasegaran et al. (2024)) proposes a benchmark dataset specifically designed for hour-long video-language understanding.
Traditional TAMP’s domain representations are usually manually specified by expert users such as PDDL (Silver et al. (2022)). However, LLMs have been explored for processing and interpreting natural language inputs (Huang et al. (2022)). They offer a novel approach to encoding the planning domain in a more intuitive and accessible way. Furthermore, LLMs’ acquisition of world knowledge and commonsense reasoning has the potential to improve the scalability and generalizability of skill hierarchy tasks (Vemprala et al. (2023); Jansen (2020); Driess et al. (2023)). Various benchmarks such as PlanBench (Valmeekam et al. (2023)) can assess the planning and reasoning capability of LLMs.
LLMs possess a notable limitation: they lack practical experience, hindering their utility for decision-making within a specific context, so the output of LLMs often cannot be translated into executable actions for the robot. Huang et al. (2022) first use pre-trained causal LLM to break down high-level tasks into logical mid-level action plans. Then, a pre-trained masked LLM is employed to convert mid-level action plans into admissible actions. However, prompts usually require the context of the robot’s capability, its current state, and the environment. LLMs are considered ”forgetful” and don’t treat information in the system prompt as absolute. Despite efforts to reinforce task constraints in the objective prompt and extract numerical task contexts from the system prompt, storing them in data structures, errors caused by LLM forgetfulness remain unresolved (Chen and Huang (2023)).
@@
Regarding user preferences, TidyBot (Wu et al. (2023b)) trains LLM by collecting users’ preference data, enabling the trained LLM to choose behaviors that better align with user preferences. As for safety regulations, Yang et al. (2023d) incorporate ISO 61508, a global standard for safely deploying robots in industrial factory settings, into the constraints of the action sequence generation. As for the skill library, BOSS (Zhang et al. (2023c)) suggests using LLMs’ rich knowledge to guide skills chaining in the skill library, aiming to create new skills through combinations. RoboGen (Wang et al. (2023e)) employs generative models to create new skill task scenarios, then utilizes either reinforcement learning or gradient optimization methods to automatically learn new skills based on the reward function generated by the LLM. As for action execution feedback, REACT (Yao et al. (2022b)), COWP (Ding et al. (2022)), LLM-Planner (Song et al. (2023)), CoPAL (Joublin et al. (2023)) and PROGPROMPT (Singh et al. (2023)) provide feedback on robot action execution to LLMs. This allows LLMs to adjust action sequences based on execution status, creating a closed-loop process for generating action sequences.
@@
Regardless of whether the prompting input to LLMs is in natural language or PDDL format, the hierarchy of skills possessed by LLMs still exhibits instability (Silver et al. (2022)). Hence, researchers are exploring approaches that integrate LLMs with classical PDDL-based planning methods for the hierarchy of skills. LLM+P (Liu et al. (2023a)) utilize LLMs to translate natural language into PDDL and input into a classical planner for the hierarchy of skills. Xie et al. (2023b) indicate that LLMs exhibit greater efficacy in translation tasks as opposed to planning.
@@
According to Fig. 5, the hierarchy of skills is mainly divided into video instruction and language instruction. VFM and LLM play roles in perception and reasoning. Language instruction is further divided into methods based on foundation models and methods combining foundation models with classical TAMP. As shown in Appendix.C Tab. 6, there is currently no significant research comparing video instruction and language instruction. However, from the modality perspective, video provides more temporal or spatial dependencies regarding tasks compared to language. This also means that video instruction requires a higher level of hierarchy of skills, not only needing to output task plans but also understanding the task and scene constraints from the video. Language instruction is more suitable for interaction and reasoning for LLMs/VLMs. The two share some similarities. Current research on the hierarchy of skills in language and video instruction tends to focus on SOTA VLMs, and both have similar failure modes, indicating that both face challenges in perception and reasoning. In the language instruction methods, the combination of foundation models and classical TAMP is more reliable than foundation models alone, but it also faces limitations in generalization. Therefore, how to better integrate foundation models with classical TAMP requires further research.
@@ 6 State
These consumer cameras typically use principles such as structured light, time of flight, and traditional photometric stereo for depth estimation (Zhou et al. (2024a)) and 3D representation can be generated by registering them using camera poses (Huang et al. (2024b)). However, when surfaces are shiny, bright, transparent, textureless or distant from the camera, depth images produced by consumer cameras are often noisy and incomplete. Several studies have addressed this challenging problem by learning to restore depth images (Dai et al. (2022); Fang et al. (2022); Sajjan et al. (2020)). However, the correct depth information may already be lost in the original depth. ASGrasp (Shi et al. (2024)) demonstrates that 3D reconstruction using raw multi-view images from consumer cameras is better than restoring the original depth. Many studies use a single image for 3D reconstruction (Fu et al. (2021)). However, a single image loses a significant amount of information, resulting in lower accuracy. Despite this, the zero-shot capability of current single image 3D reconstruction has led to its widespread application in simulation scene generation (in Sec. 8.3).
The 3D representation for 3D reconstruction can be expressed as explicit and implicit expressions (Zhou et al. (2024a)). Explicit expressions include point clouds (Shi et al. (2024)), voxels (Jiang et al. (2021)), and meshes (Wen et al. (2019)). The three representations can be converted into each other (Jiang et al. (2021)), but each has its own advantages. A point cloud is made up of discrete points in space, providing flexibility in processing. A voxel can store spatial information inside an object but comes with high space complexity. A mesh uses triangle meshes to represent complex shapes and details accurately, such as deformation (Wen et al. (2019)). It ensures the projection is always convex, making it easier to rasterize (Zhou et al. (2024a)). Implicit expressions represent 3D geometry using a function, such as signed distance function (SDF) (Chabra et al. (2020)), occupancy field (Jiang et al. (2021)), and radiance field (Mildenhall et al. (2021)). They offer differentiability and efficient storage, making them a powerful tool. In contrast, explicit expressions tend to be more intuitive.
@@
The image features output by NeRF can be used to build a relevancy map. They can also be lifted into 3D space through multi-view images, serving as 3D features for downstream tasks (Ze et al. (2023)). 3D-LLM (Hong et al. (2023b)) extracts 2D features from multi-view rendered images using the CLIP image encoder. These features are then fused into 3D features through Direct Reconstruct, gradSLAM (Jatavallabhula et al. (2023)), or Neural Field methods (Hong et al. (2023b)), endowing 3D features with semantic information.